Explain how the production-possibilities frootier of the unified Germany might differ from the PPF of the former Federal Republic of Germany (West Germany),keeping in mind that West Germany, in the two-factor context, was generally considered relatively capital abundant and the German Democratic Republic (East Germany) was generally considered relatively labor abundant. What would theory suggest about the differences in relative output of capital intensive goods and labor-intensive goods of the former West Germany compared with the unified Germany? What would theory suggest, if anything. about the rade pattern of the new Gemany compared with that of the former West Germany if it is assumed that the former West Germany was capital abuadant relative to its trading partners?
6. The nominal tariff rates on the 10 imports into the fictional country of Tarheelia, as well as the total import value of each good, are listed bere Valwe $400 S600 S500 300 S200 S400 100 400 (100 units $200 $100 (10 units) Nominal Rate 10% 5% Free 30% Good A Good B Good C Good D Good E Good F Good G Good H Good I Good J 2.5% 15% S0.50 unit 40% $2.50/unit (a) Calculate the unweighted-average nominal tariff rate for Tarheelia. (b) Calculate the weighted-average nominal tariff rate for Tarheelia.
Answer #1

Answer:

Question 6)

Nominal rate of Good C=0%

Nominal rate of Good H=0.5*100/400=12.5%

Nominal rate of Good J=2.5*10/100=25%

Goods Nominal Rate Nominal Rate Ni Value Vi
A 10% 0.1 400
B 5% 0.05 600
C free 0% 500
D 30% 0.3 300
E 2% 0.02 200
F 2.50% 0.025 400
G 15% 0.15 100
H $0.5/unit 12.50% 400
I 40% 0.4 200
J $2.5/unit 25% 100

A) unweighted average nominal rate =Ni/10

So unweighted average nominal rate=142%/10=14.2%

B) weighted average nominal rate =(sum Vi*Ni)/(sum Vi)

So unweighted average nominal rate=344/3200=10.75%